OpenStudy (anonymous):

Solve 5x = 50. Round your answer to the nearest hundredth. Answer x ≈ 2.43 x ≈ 4.56 x ≈ 3.89 x ≈ 1.67

OpenStudy (mathstudent55):

You mean \(5^x = 50\), right? Are we supposed to read your mind?

OpenStudy (anonymous):

Lolz. I think thats what he meant, too!

OpenStudy (anonymous):

um...5x = 50.....x = 10....is your problem written wrong ?

OpenStudy (anonymous):

no thats what the question asks me & no i know you cant read my mind !

OpenStudy (mathstudent55):

\(5^x = 50\) Take log of both sides: \(\log 5^x = \log 50\) \(x \log 5 = \log 50 \) \(x = \dfrac{\log 50}{\log 5} \)

OpenStudy (anonymous):

gotta love those exponents !

OpenStudy (anonymous):

If the problem is 5^x = 50 x = (log(2)+2 log(5))/(log(5)) x = 2.4307 http://www.wolframalpha.com/input/?i=5%5Ex+%3D+50

OpenStudy (mathstudent55):

You wrote 5x = 50, but your answers are to \(5^x = 50\). If we solve the problem you actually asked for, the answer is x = 10. If, by reading your mind, we answer the problem you meant to ask, then the answer is \(x = \dfrac{\log 50}{\log 5} \approx 2.43\)

OpenStudy (anonymous):

thanks math students &ilfy

OpenStudy (anonymous):

No problem!

OpenStudy (mathstudent55):

You're welcome. In the future, could you please use the exponent symbol ^ to make it easier to interpret your question correctly? All you need to do was write: 5^x = 50